OpenStudy (anonymous):

Your heart rate is the number of heartbeats per minute. What is your heart rate if you count 18 beats in 15 seconds? How many beats do you count in 15 seconds if your heart rate is 96 beats/minute?

OpenStudy (smileyxl3):

60/15=4 18*4=72 = heart rate --------------------- 96/4=24 beats per 15 seconds. Did you get that?

OpenStudy (anonymous):

Yes, I see thank you